# A Context for Manifold Calculus

### v2

@@ A.1. A Remark on Classifying Spaces of Groupoids Internal to $\mathsf{Top}$
  • In [McD80], McDuff defines the classifying space of $\mathcal{G}$ to be the fat realization of the simplicial topological space $\mathcal{G}$. We denote this topological space by $B_{{\rm MD}}\left(\mathcal{G}\right)$ [^5] This model was introduced by Segal [Seg68]..
@@
    Now introduce an equivalence relation on $E_{{\rm Hae}}\mathcal{G}$ as follows: Two points $\bigoplus_{j\geq 0}t_{j}g_{j}$ and $\bigoplus_{j\geq 0}t^{\prime}_{j}g^{\prime}_{j}$ are equivalent if there is some morphism $g$ of $\mathcal{G}$ such that $\bigoplus_{j\geq 0}t_{j}gg_{j}=\bigoplus_{j\geq 0}t_{j}gg^{\prime}_{j}$. The resulting quotient space will be denoted by $B_{{\rm Hae}}\mathcal{G}$. We still denote by $\tau_{i}:B_{{\rm Hae}}\mathcal{G}\to[0,1]$ the map induced by $\tau_{i}:E_{{\rm Hae}}\mathcal{G}\to[0,1]$. The space $B_{{\rm Hae}}\mathcal{G}$ is the model of the classifying space of $\mathcal{G}$ used by Haefliger in [Hae71] [^6] This model was introduced by Milnor in [Mil56]..
@@
The proof of Proposition A.2 relies on a lemma. Write $N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)$ for the semisimplicial set of nondegenerate simplices of the nerve of the poset $\mathbb{Z}_{\geq 0}$.
@@
where $\left(t_{0},\dots,t_{n}\right)$ is a point of $\left|\Delta^{n}\right|$, $x_{0}\xrightarrow{f_{1}}\cdots\xrightarrow{f_{n}}x_{n}$ is a point of $\left(N\mathcal{G}\right)_{n}$, and $k_{0}<\cdots<k_{n}$ is an element of $N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)_{n}$. We declare that $\phi$ carries such a point to the point
@@
Proof of A.2.
@@
where $\phi$ is the map of Lemma A.3 and $\phi^{\prime}$ is induced by the projection. We claim that these maps are weak homotopy equivalences. The map $\phi$ is a weak homotopy equivalence by Lemma A.3. For the map $\phi^{\prime}$, notice that the map

### v3

@@ A.3. A Remark on Classifying Spaces of Groupoids Internal to $\mathsf{Top}$
  • In [McD80], McDuff defines the classifying space of $\mathcal{G}$ to be the fat realization of the simplicial topological space $\mathcal{G}$. We denote this topological space by $B_{{\rm MD}}\left(\mathcal{G}\right)$. [^6] This model was introduced by Segal [Seg68].
@@
    Now introduce an equivalence relation on $E_{{\rm Hae}}\mathcal{G}$ as follows: Two points $\bigoplus_{j\geq 0}t_{j}g_{j}$ and $\bigoplus_{j\geq 0}t^{\prime}_{j}g^{\prime}_{j}$ are equivalent if there is some morphism $g$ of $\mathcal{G}$ such that $\bigoplus_{j\geq 0}t_{j}gg_{j}=\bigoplus_{j\geq 0}t_{j}gg^{\prime}_{j}$. The resulting quotient space will be denoted by $B_{{\rm Hae}}\mathcal{G}$. We still denote by $\tau_{i}:B_{{\rm Hae}}\mathcal{G}\to[0,1]$ the map induced by $\tau_{i}:E_{{\rm Hae}}\mathcal{G}\to[0,1]$. The space $B_{{\rm Hae}}\mathcal{G}$ is the model of the classifying space of $\mathcal{G}$ used by Haefliger in [Hae71]. [^7] This model was introduced by Milnor in [Mil56].
@@
[…]
@@
where $\left(t_{0},\dots,t_{n}\right)$ is a point of $\left|\Delta^{n}\right|$, $x_{0}\xrightarrow{f_{1}}\cdots\xrightarrow{f_{n}}x_{n}$ is a point of $\mathcal{G}_{n}$, and $k_{0}<\cdots<k_{n}$ is an element of $N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)_{n}$. We declare that $\phi$ carries such a point to the point
@@
For the next lemma, we write $\mathbf{\Delta}_{\mathrm{inj}}\subset\mathbf{\Delta}$ for the subcategory spanned by the injective poset maps. We also identify sets with discrete simplicial sets. Note that homotopy colimits over $\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}$ do not commute with products (up to weak equivalence) in general, as its nerve is not sifted.
Lemma A.15.
Let $\mathcal{C}$ be a category internal to $S\SS$. The map
$\operatorname{hocolim}_{\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\left(\mathcal{C}\times N^{\mathrm{nd}}\left(\mathbb{Z}_{\geq 0}\right)\right)\to\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\mathcal{C}$
is a weak homotopy equivalence.
Remark A.16.
Note that in general homotopy colimits over $\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}$ do not commute with products even up to weak equivalence, as its nerve is not sifted. In contrast, the nerve of $\mathbf{\Delta}^{\mathrm{op}}$ is sifted [Lur09, Lemma 5.5.8.4].
Proof.
[…]
$\Phi\left(A\right)=\int^{[n]\in\mathbf{\Delta}_{\mathrm{inj}}}A_{n}\times\Delta^{n}.$
To see that $SS\Phi:\SS^{\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\to\SS$ is actually a model of homotopy colimits, observe that its right adjoint $S\mapsto S^{\Delta^{\bullet}}$ is right Quillen for the Reedy model structure by inspection, and that it is naturally weakly equivalent to the diagonal functor. Since every semisimplicial simplicial set is Reedy cofibrant, this means that $\Phi$ is a model of homotopy colimits.
[…]
$\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\left(X_{n}\right)$ $\cong\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\operatorname{hocolim}_{[m]\in\mathbf{\Delta}^{\mathrm{op}}}\left(X_{n,m}\right)$
$\simeq\operatorname{hocolim}_{[m]\in\mathbf{\Delta}^{\mathrm{op}}}\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}\left(X_{n,m}\right),$
which is natural in $X$. This implies that, given a map $f:X\to Y$ of semisimplicial simplicial sets, if the map
$\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}X_{n,m}\to\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}Y_{n,m}$
is a weak homotopy equivalence for every $m\geq 0$, then $\operatorname{hocolim}_{[n]\in\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}f$ is a weak homotopy equivalence. Substituting $\mathcal{C}\times N^{\mathrm{nd}}\left(\mathbb{Z}_{\geq 0}\right)$ for $X$ and $\mathcal{C}$ for $Y$, we are reduced to the case where each $\mathcal{C}_{n}$ is a (discrete simplicial) set, i.e., $\mathcal{C}=N\left(\mathsf{C}\right)$ for some ordinary category $\mathsf{C}$.
Let $\mathsf{C}^{\prime}\subset\mathsf{C}\times\mathbb{Z}_{\geq 0}$ denote the subcategory generated by the morphisms $\left(X,n\right)\to\left(Y,m\right)$ such that $n<m$. We can identify the semisimplicial set $N\left(\mathsf{C}\right)\times N^{{\rm nd}}\left(\mathbb{Z}_{\geq 0}\right)$ with the restriction of the nerve $N\left(\mathsf{C}^{\prime}\right)$ to $\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}$. Under this identification, our goal is to show that the projection $\pi:\mathsf{C}^{\prime}\to\mathsf{C}$ gives a weak homotopy equivalence
$\operatorname{hocolim}_{\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}N\left(\mathsf{C}^{\prime}\right)\to\operatorname{hocolim}_{\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}}N\left(\mathsf{C}\right).$
Since the inclusion $\mathbf{\Delta}_{\mathrm{inj}}^{\mathrm{op}}\hookrightarrow\mathbf{\Delta}^{\mathrm{op}}$ is homotopy final (being a right adjoint), we may replace the indexing category of homotopy colimits by $\mathbf{\Delta}^{\mathrm{op}}$. With our model of $\operatorname{hocolim}_{\mathbf{\Delta}^{\mathrm{op}}}$, the relevant map will be just the nerve of $\pi$. Thus, we are reduced to showing that $N\left(\pi\right)$ is a weak homotopy equivalence.
To prove that $N\left(\pi\right)$ is a weak homotopy equivalence, we factor $\pi$ as $\mathsf{C}^{\prime}\xrightarrow{\iota}\mathsf{C}\times\mathbb{Z}_{\geq 0}\xrightarrow{\pi^{\prime}}\mathsf{C}$, where $\iota$ is the inclusion and $\pi^{\prime}$ is the projection. Since $N\left(\pi^{\prime}\right)$ is a weak homotopy equivalence (as $N\left(\mathbb{Z}_{\geq 0}\right)$ is weakly contractible), it will suffice to show that $N\left(\iota\right)$ is a weak homotopy equivalence. We prove this by using Quillen’s Theorem A: We must show that, for each $\left(X,n\right)\in\mathsf{C}\times\mathbb{Z}_{\geq 0}$, the fiber product
$\mathsf{P}=\mathsf{C}^{\prime}\times_{\mathsf{C}\times\mathbb{Z}_{\geq 0}}\left(\mathsf{C}\times\mathbb{Z}_{\geq 0}\right)_{\left(X,n\right)/}$
has weakly contractible nerve. The inclusion $\left(\mathsf{C}^{\prime}\right)_{\left(X,n\right)/}\hookrightarrow\mathsf{P}$ is a left adjoint (with right adjoint given by $\left(Y,m\right)\mapsto\left(X,n\right)$ if $m=n$ and $\left(Y,m\right)\mapsto\left(Y,m\right)$ if $m>n$), so it suffices to show that $N\left(\left(\mathsf{C}^{\prime}\right)_{\left(X,n\right)/}\right)$ is weakly contractible. This is clear, as $\left(\mathsf{C}^{\prime}\right)_{\left(X,n\right)}$ has an initial object. The proof is now complete. ∎
Proof of Proposition A.13.
@@
where $\phi$ is the map of Lemma A.14 and $\phi^{\prime}$ is induced by the projection. The map $\phi$ is a weak homotopy equivalence by Lemma A.14. We will complete the proof by showing that $\phi^{\prime}$ is a weak homotopy equivalence.
[…]
In light of the discussion in the previous paragraph, we can identify $\phi^{\prime}$ with the map